Paul and Donna Decker are married taxpayers, ages 44 and 42, respectively, who file a joint return for 2019. The Deckers live at 1121 College Avenue, Carmel, IN 46032. Paul is an assistant manager at Carmel Motor Inn, and Donna is a teacher at Carmel Elementary School. They present you with W–2 forms that reflect the following information:

Paul Donna
Salary $68,000 $56,000
Federal tax withheld 6,770
6,630
State income tax withheld 1,400 1,100
FICA (Social Security and Medicare) withheld 5,202 4,284
Social Security numbers 111-11-1112
123-45-6789

Donna is the custodial parent of two children from a previous marriage who reside with the Deckers throughout the school year. The children, Larry and Jane Parker, reside with their father, Bob, during the summer. Relevant information for the children follows:

Larry Jane
Age      17 18
Social Security numbers 123-45-6788 123-45-6787
Months spent with Deckers         9 9

Under the divorce decree, Bob pays child support of $150 per month per child during the nine months the children live with the Deckers. Bob says that he spends $200 per month per child during the three summer months they reside with him. Donna and Paul can document that they provide $2,000 support per child per year. The divorce decree is silent as to which parent can claim the exemptions for the children.

In August, Paul and Donna added a suite to their home to provide more comfortable accommodations for Hannah Snyder (123-45-6786), Donna’s mother, who had moved in with them in February 2018 after the death of Donna’s father. Not wanting to borrow money for this addition, Paul sold 300 shares of Acme Corporation stock for $50 per share on May 3, 2019, and used the proceeds of $15,000 to cover construction costs. The Deckers had purchased the stock on April 29, 2014, for $25 per share. They received dividends of $750 on the jointly owned stock a month before the sale.

Hannah, who is 66 years old, received $7,500 in Social Security benefits during the year, of which she gave the Deckers $2,000 to use toward household expenses and deposited the remainder in her personal savings account. The Deckers determine that they have spent $2,500 of their own money for food, clothing, medical expenses, and other items for Hannah. They do not know what the rental value of Hannah’s suite would be, but they estimate it would be at least $300 per month.

Interest paid during the year included the following:

Home mortgage interest (paid to Carmel Federal Savings & Loan) $7,890
Interest on an automobile loan (paid to Carmel National Bank) 1,660
Interest on Citibank Visa card 620

In July, Paul hit a submerged rock while boating. Fortunately, he was uninjured after being thrown from the boat and landing in deep water. However, the boat, which was uninsured, was destroyed. Paul had paid $25,000 for the boat in June 2018, and its value was appraised at $18,000 on the date of the accident.

The Deckers paid doctor and hospital bills of $12,700 and were reimbursed $2,000 by their insurance company. They spent $640 for prescription drugs and medicines and $5,904 for premiums on their health insurance policy. They have filed additional claims of $1,200 with their insurance company and have been told they will receive payment for that amount in January 2020. Included in the amounts paid for doctor and hospital bills were payments of $380 for Hannah and $850 for the children.

Additional information of potential tax consequence follows:

Real estate taxes paid $6,850
Sales taxes paid (per table) 1,379
Contributions to church 4,600
Appraised value of books donated to public library 740
Refund of state income tax for 2018 (the Deckers itemized on their 2018 Federal tax return, and their total state and local taxes were less than $10,000) 1,520

Compute net tax payable or refund due for the Deckers for 2019. If the Deckers have overpaid, the amount is to be credited toward their taxes for 2020.
Answer #1

1040 and

First name : Paul

LastName: Decker

SSN: 111-11-1111

Spousefirst name:Donna

LastName:Decker

SSN:123-45-6789

Home address :1121 College Avenue,

City:Carmel state:IN   ZIPcode:46032.

Filing Status : 2)Married filing jointly

Exemption

6C)First name: Last Name:   Social securitynumber: Relationship: Child age

Larry Decker 123-45-6788 Son 17

Jane Decker 123-45-6787 Daughter 18

Hannah Snyder 123-45-6786 Mother in law N/A

6d) Total number of exemptions: 3

Income:

7)Wages: 68,000 + 56,000=124,000

9 a) Ordinary dividends.$750

10) Taxable refunds: $1,520

13) Capital gain: $7,500

Acme Corporation stock: 300 shares * $50=$15,000 - 300*$25 =$7500

Gain = $15,000 -$7500=$7,500:

22) Combine the amounts in the far right column for lines 7through 21. $133,770

37) Subtract line 36 from line 22: $133,770

This is your adjusted gross income

38)Amount from line 37: $133,770

40) Itemized deductions (from Schedule A) : $43,836

41) Subtract line 40 from line 38: $133,770 -$43,836 =$89934

42) Exemptions: 6D 3 * 4050=$12,150

43) Taxable income: $89934 -$12,150=$77784

Subtract line 42 from line 41.

64) Federal income tax withheld from =6,770 +6,630=$13,400

65) 2016 estimated tax payments: $10988.50

upto75300 =10367.50 + $621: $77784 - 75300=2484*25/100=$621

74) Add lines 64, 65, 66a, and 67 through 73.

These are your total payments: $13,400 +$10988.50=$24,388.50

1)Medical and dental expenses : $16,444

2) Enter amount from Form 1040, line 38: $133,770

3) Multiply line 2 by 10% (0.10): 13,377

4) Subtract line 3 from line 1. =$16,444 - 13,377=$3067

TAX:

5 b General sales taxes:1,379

6) Real estate taxes: 3,850

9) Add lines 5 through 8: 1,379 +3,850=$5229

INTEREST

10)Home mortgage interest:7,890

11) Home mortgage interest :1,660

15) Add lines 10 through 14:9550

GIFTS:

16 gifts by cash or check :$250(ALLOWED)

17)OTHER THAN CASH 740

19) Add lines 16 through 18: $990

20 Casualty or theft loss : $25,000

21 Unreimbursedemployee: $340+$340+$170+$47.50=$897.50

24) Add lines 21 through 23: $897.50

25) Enter amount from Form 1040, line 38: $133,770

26) Multiply line 25 by 2%(0.02): $133,770 *2/100=$2675.40

27) Subtract line 26 from line 24: 0

29) Total Itemized Deductions : $3067 + $5229 +9550 + $990+$25000=$43,836.

MEDICAL CALCULATIONS:

Hospital bills = $10,700 +$1,200 +$640 +$5,904 -$2,000 =$16,444